OpenStudy (anonymous):

the value of a car declines 20% per year. what percent of the original value remains after 2 years?

OpenStudy (oaktree):

If the car is worth $1,000...

OpenStudy (oaktree):

After the first year, then it'll be worth $800.

OpenStudy (oaktree):

Then, it'll be worth 80/100 x 800 = 80 x 8 = $640.

OpenStudy (anonymous):

what if it is worth $100

OpenStudy (oaktree):

So, 640/1000 = 64/100 = 64%.

OpenStudy (oaktree):

It's the same no matter what. Because every time you're reducing by 20%.

OpenStudy (oaktree):

If the car is worth x, for example, then after the first year the car is 80%x and after the second it's worth 80%(80%x) = 4/5 x 4/5 = 16/25 = 64/100 = 64%.

OpenStudy (oaktree):

Got it?

OpenStudy (anonymous):

yes!! thank you so much

OpenStudy (anonymous):

great job OakTree!